OpenStudy (anonymous):

Multitply (x^8+2x)(x^8-2x)=

OpenStudy (anonymous):

x^16 - 4x^2

jimthompson5910 (jim_thompson5910):

(x^8+2x)(x^8-2x) (x^8)(x^8)+(x^8)(-2x)+(2x)(x^8)+(2x)(-2x) x^16-2x^9+2x^9-4x^2 x^16-4x^2 So (x^8+2x)(x^8-2x)=x^16-4x^2

OpenStudy (anonymous):

(a-b)(a+b) = a^2-b^2 \[(x^8+2x)(x^8-2x)\] that would be (x^8)^2 - (2x)^2 x^16 - 4x^2 \[x ^{16}-4x^{2} \]